The Daytona 500 started in 1959 with an average speed of 135 mph but before they raced on the big speedway they would race up and down on the beach. In 1954 Lee Petty raced on the beach for a total 160 miles. As the 2 hour race developed Lee averaged 90 mph for the first part of the race but only 60 mph for last part. How long did he race at 90 mph and how long at 60 mph?
Answer by richwmiller(17219)

x+y=2
x=2-y
90(2-y)+60y=160
20=30y
y=2/3 hr at 60 mph
x=4/3 hr at 90 mph
check
90*4/3+60*2/3
120+40=160
ok